OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (5, 9); Slope: 2

OpenStudy (anonymous):

A. y – 9 = 2(x + 5) B. y – 5 = 2(x + 9) C. y + 9 = 2(x – 5) D. y – 9 = 2(x – 5)

OpenStudy (anonymous):

its c or a

undeadknight26 (undeadknight26):

Looks like C...i think it would only make it a addition sign if there was a negative number as x1.

OpenStudy (anonymous):

thanks